# UniAP: Unifying Inter- and Intra-Layer Automatic Parallelism by Mixed Integer Quadratic Programming

### v4

@@ 1 Introduction
Distributed learning (also called parallel learning) on clusters with several machines or GPUs is commonly used for training deep learning models, especially for some large models with billions of parameters [2; 6; 7]. Several parallel strategies, including pipeline parallelism (PP), data parallelism (DP), tensor parallelism (TP), and fully sharded data parallelism (FSDP), have been proposed for distributed learning. These parallel strategies can be divided into two main categories: inter-layer parallelism and intra-layer parallelism. Inter-layer parallelism [8; 9; 10; 11; 12; 13; 14; 15], which includes PP, partitions the model into disjoint sets without partitioning tensors in each layer. Intra-layer parallelism [16; 17; 18; 19], which includes DP, TP, and FSDP, partitions tensors in a layer along one or more axes.
The parallel method [^1] To avoid confusion, we treat ‘parallel method’ and ‘parallel strategy’ as two different terminologies in this paper. in one specific distributed learning method or system typically adopts one parallel strategy or a combination of several parallel strategies. Existing parallel methods can be divided into two categories: manual parallelism (MP) methods and automatic parallelism (AP) methods. In MP methods  [20; 18; 21], one or several parallel strategies are manually optimized by researchers or developers. MP methods require extensive domain knowledge in deep learning models and hardware architectures. With the rapid development of deep learning models and the increasing diversity of modern hardware architectures [22; 23], MP methods demand considerable human effort and have limited flexibility.
To address the two limitations of MP methods, AP methods [9; 24; 25] have recently been proposed for automating the parallel strategy optimization process. Although existing AP methods have achieved promising progress, they optimize the two categories of parallel strategies separately rather than jointly. More specifically, some methods optimize only one category of parallel strategies [26; 27; 28; 29; 30; 31; 32], and the others optimize inter- and intra-layer parallelism hierarchically [9; 33; 34; 11; 24; 25]. Hence, existing AP methods suffer from sub-optimal solutions.
@@ 3 Method
In this section, we introduce our proposed method called UniAP. Apart from current AP methods, UniAP jointly optimizes the two categories of parallel strategies, including PP, DP, TP, and FSDP, to find an optimal solution. Visualization on the differences between existing parallel methods and UniAP is available in Appendix A.
@@ 3.1 Profiling
UniAP collects runtime information about the hardware environment and deep learning model during profiling. For the hardware environment, UniAP evaluates the efficiency of all-reduce and point-to-point (P2P) communication for different device subsets. For example, when profiling a node with 4 GPUs, UniAP measures the all-reduce efficiency for various DP, TP, and FSDP combinations across these GPUs. Additionally, UniAP ranks these GPUs from 0 to 3 and evaluates the speed of P2P for two pipeline options: ($0\rightarrow 2$ and $1\rightarrow 3$) and ($0\rightarrow 1$, $1\rightarrow 2$ and $2\rightarrow 3$). Furthermore, UniAP estimates the computation-communication overlap coefficient (CCOC) [37; 38].
@@ 3.2 Cost model
Time cost model To estimate computation time, UniAP first calculates the forward computation time by multiplying the batch size with the forward computation time per sample obtained from profiling. Users can obtain a more precise result by profiling the forward time on the specified batch size. For Transformer-based models that mainly consist of the MatMul operator, the computation time in the BP stages is roughly twice that of the FP stages [18; 12; 37]. Additionally, UniAP estimates the communication time by dividing the size of transmitting tensors by the profiled communication efficiency for different communication primitives. To accommodate overlapping, UniAP multiplies the profiled CCOC by the overlapping interval of computation and communication. To model the communication time between pipeline stages, UniAP calculates the cross-stage cost between consecutive stages by the summation of P2P costs.
@@
For example, if we choose precision as FP32, then $c_{dtype}=(4+4+4+4)/4=4$ since the learnable parameters, gradients, momentum and variance consume equal size of memory. If we opt for mixed precision with FP16 activated, then $c_{dtype}=(4+4+4+2+2)/2=8$. Finally, UniAP aggregates the activation memory cost $m_{a}$, memory cost of model states $m_{s}$, and context memory cost $m_{c}$ to a constant matrix ${\boldsymbol{\mathit{M}}}$, where ${\boldsymbol{\mathit{M}}}_{uk}$ denotes the memory cost for the $k$-th intra-layer strategy of layer $u$ on a single device.
@@ 3.3.1 Objective function
The objective function tries to minimize TPI. In this paper, we have chosen GPipe as our PP strategy for illustration. [^2] UniAP is also compatible with other PP strategies. For example, users need to modify only the memory constraint in Section 3.3.2 to adapt to synchronous 1F1B pipeline [10; 11]. Figure 3 depicts the time cost decomposition of a GPipe-style PP with non-negligible communication costs. The time needed to apply gradients at the end of each iteration is not included, as it depends on the optimizer and is insignificant compared to the total time spent on FP and BP.
We denote the cost for computation stages as =$\mathbb{P}=\{p_{1},p_{2},\dots,p_{pp\_size}\}$ and the cost for communication stages as $\mathbb{O}=\{o_{1},o_{2},\dots,o_{pp\_size-1}\}$. Here, $pp\_size$ represents the number of computation stages, which corresponds to the pipeline parallel size. In Figure 3, $fp_{i}$ and $bp_{i}$ denote forward and backward computation time for computation stage $i$, respectively. $fo_{j}$ and $bo_{j}$ denote forward and backward communication time for communication stage $j$, respectively. Hence, we have $p_{i}=fp_{i}+bp_{i}$ and $o_{j}=fo_{j}+bo_{j}$.
@@
$\min tpi_{gpipe}=\sum_{i=1}^{pp\_size}p_{i}+\sum_{j=1}^{pp\_size-1}o_{j}+(c-1)\max\left(\mathbb{P}\cup\mathbb{O}\right).$ (2)
@@ 3.3.2 Constraint
We first introduce additional notations before presenting the constraints. For a given layer $u\in{\mathbb{V}}$, ${\mathbb{S}}_{u}$ represents its set of intra-layer parallel strategies, ${\boldsymbol{\mathit{A}}}_{uk}$ denotes the $k$-th intra-layer execution cost obtained from our time cost model. Additionally, we use ${\boldsymbol{\mathit{S}}}_{uk}\in\{0,1\}$ to indicate whether the $k$-th parallel strategy is selected for the layer $u$, and use ${\boldsymbol{\mathit{P}}}_{ui}\in\{0,1\}$ to indicate whether layer $u$ is placed on the $i$-th computation stage. Each edge $\langle u,v\rangle\in{\mathbb{E}}$ is assigned a resharding cost denoted by ${\boldsymbol{\mathit{R}}}_{uv}$ if the vertices are located within the same pipeline stage. Alternatively, if the vertices are located across consecutive stages, the resharding cost between them is denoted by ${\boldsymbol{\mathit{R}}}^{\prime}_{uv}$. These two resharding costs are constant matrices derived from our time cost model.
@@
$\sum_{u\in{\mathbb{V}}}{\boldsymbol{\mathit{P}}}_{ui}{\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{A}}}_{u}+\sum_{\langle u,v\rangle\in{\mathbb{E}}}{\boldsymbol{\mathit{P}}}_{ui}{\boldsymbol{\mathit{P}}}_{vi}({\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{R}}}_{uv}{\boldsymbol{\mathit{S}}}_{v})=p_{i},~{}\forall i\in\{1,\dots,pp\_size\}.$ (3)
@@
$\sum_{\langle u,v\rangle\in{\mathbb{E}}}{\boldsymbol{\mathit{P}}}_{uj}{\boldsymbol{\mathit{P}}}_{v(j+1)}({\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{R}}}^{\prime}_{uv}{\boldsymbol{\mathit{S}}}_{v})$ $=o_{j},~{}\forall j\in\{1,\dots,pp\_size-1\}.$ (4)
@@
$\sum_{u\in{\mathbb{V}}}{\boldsymbol{\mathit{P}}}_{ui}{\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{M}}}_{u}\leqslant m,~{}\forall i\in\{1,\dots,pp\_size\}.$ (5)
@@
Order-preserving constraint PP is not a single-program multiple-data (SPMD) parallel strategy [8]. Hence, we need an order-preserving constraint to ensure that the subgraphs of $\mathcal{G}$ are contiguous. We adopt the definition of contiguous from previous work [33; 34].
@@
${\boldsymbol{\mathit{Z}}}_{vi}$ $\geqslant{\boldsymbol{\mathit{P}}}_{vi},~{}$ $\forall v\in{\mathbb{V}},~{}\forall i\in\{1,2,\dots,pp\_size\},$ (6a)
${\boldsymbol{\mathit{Z}}}_{vi}$ $\leqslant{\boldsymbol{\mathit{Z}}}_{ui},~{}$ $\forall u,v\in{\mathbb{V}},~{}\forall\langle u,v\rangle\in{\mathbb{E}},~{}\forall i\in\{1,2,\dots,pp\_size\},$ (6b)
${\boldsymbol{\mathit{Z}}}_{vi}$ $\leqslant{\boldsymbol{\mathit{P}}}_{vi}-{\boldsymbol{\mathit{P}}}_{ui}+1,~{}$ $\forall u,v\in{\mathbb{V}},~{}\forall\langle u,v\rangle\in{\mathbb{E}},~{}\forall i\in\{1,2,\dots,pp\_size\}.$ (6c)
Detailed proof can be found in Appendix B.
@@
$\sum_{i=1}^{pp\_size}{\boldsymbol{\mathit{P}}}_{ui}=1,$ $\forall u\in{\mathbb{V}},$ (7a)
$\sum_{u\in{\mathbb{V}}}{\boldsymbol{\mathit{P}}}_{ui}\geqslant 1,$ $\forall i\in\{1,\dots,pp\_size\},$ (7b)
${\boldsymbol{\mathit{P}}}_{ui}\in\{0,1\},$ $\forall u\in{\mathbb{V}},~{}i\in\{1,\dots,pp\_size\}.$ (7c)
@@
$\sum_{k=1}^{\lvert{\mathbb{S}}_{u}\rvert}{\boldsymbol{\mathit{S}}}_{uk}=1,$ $\forall u\in{\mathbb{V}},$ (8a)
${\boldsymbol{\mathit{S}}}_{uk}\in\{0,1\},$ $\forall u\in{\mathbb{V}},~{}k\in\{1,\dots,|{\mathbb{S}}_{u}|\}.$ (8b)
@@ 3.4 Unified optimization process
  Output: Optimal cost $cost^{*}$, pipeline parallel size $pp\_size^{*}$, the number of micro-batches $c^{*}$, layer placement ${\boldsymbol{\mathit{P}}}^{*}$, and intra-layer strategy ${\boldsymbol{\mathit{S}}}^{*}$
  $pp\_size^{*}=1$;
@@
  $cost^{*}$, $P^{*}$, $S^{*}$ = QIP(${\boldsymbol{\mathit{A}}}$, ${\boldsymbol{\mathit{R}}}$, ${\boldsymbol{\mathit{M}}}$);
@@
        ${\boldsymbol{\mathit{A}}}$, ${\boldsymbol{\mathit{R}}}$, ${\boldsymbol{\mathit{R}}}^{\prime}$, ${\boldsymbol{\mathit{M}}}$ = CostModeling($PR$, $SD[pp\_size]$, $\mathcal{G}$, $b$);
        $cost$, ${\boldsymbol{\mathit{P}}}$, ${\boldsymbol{\mathit{S}}}$ = MIQP(${\boldsymbol{\mathit{A}}}$, ${\boldsymbol{\mathit{R}}}$, ${\boldsymbol{\mathit{R}}}^{\prime}$, ${\boldsymbol{\mathit{M}}}$, $pp\_size$, $c$);
@@
           $cost^{*}$, $pp\_size^{*}$, $c^{*}$, ${\boldsymbol{\mathit{P}}}^{*}$, ${\boldsymbol{\mathit{S}}}^{*}$ = $cost$, $pp\_size$, $c$, ${\boldsymbol{\mathit{P}}}$, ${\boldsymbol{\mathit{S}}}$;
@@
UOP integrates the cost model and MIQP based on the profiling results and the computation graph to return the optimal parallel strategy and the corresponding TPI. Algorithm 1 summarizes the whole process. In Algorithm 1, we denote intra-layer cost as ${\boldsymbol{\mathit{A}}}$, inter-layer cost as ${\boldsymbol{\mathit{R}}}$, cross-stage cost as ${\boldsymbol{\mathit{R}}}^{\prime}$, and memory cost as ${\boldsymbol{\mathit{M}}}$. The CostModeling process calculates these four costs based on the cost model described in Section 3.2.
First, UOP optimizes intra-layer-only parallelism for cases in which pipeline parallelism is not adopted. Several works [25; 32] have used quadratic integer programming (QIP) to optimize intra-layer-only parallel strategy and achieved promising results. UniAP provides a QIP formulation for intra-layer-only parallelism in Appendix C.
Then, UOP enumerates all factors of $n$ except 1 as the pipeline parallel size $pp\_size$. For each $pp\_size$, UOP enumerates all factors of $B$ except 1 as the number of micro-batches $c$. These enumerations aim to achieve load balance on a homogeneous cluster.
For each candidate $pp\_size$ and $c$, UOP formulates the cost for a training iteration to an MIQP expression. It then waits for the MIQP solver to return the optimal cost and parallel strategy under the current configuration. Specifically, our implementation adopts the Gurobi Optimizer [39] as our QIP and MIQP solver.
Finally, UOP returns the minimum cost $cost^{*}$ and its corresponding pipeline parallel size $pp\_size^{*}$, number of micro-batches $c^{*}$, layer placement ${\boldsymbol{\mathit{P}}}^{*}$, and intra-layer strategies ${\boldsymbol{\mathit{S}}}^{*}$. We provide visualization for a candidate solution to UOP in Appendix D.

### v5

@@ 1 Introduction
Distributed learning (also called parallel learning) on clusters with several machines or GPUs is commonly used for training deep learning models, especially for some large models with billions of parameters [1, 40, 41]. Several parallel strategies, including pipeline parallelism (PP), data parallelism (DP), tensor parallelism (TP), and fully sharded data parallelism (FSDP), have been proposed for distributed learning. These parallel strategies can be divided into two main categories: inter-layer parallelism and intra-layer parallelism. Inter-layer parallelism [15, 28, 29, 9, 21, 20, 7, 10], which includes PP, partitions the model into disjoint sets without partitioning tensors in each layer. Intra-layer parallelism [22, 34, 30, 8], which includes DP, TP, and FSDP, partitions tensors in a layer along one or more axes.
The parallel method [^1] To avoid confusion, we treat ‘parallel method’ and ‘parallel strategy’ as two different terminologies in this paper. in one specific distributed learning method or system typically adopts one parallel strategy or a combination of several parallel strategies. Existing parallel methods can be divided into two categories: manual parallelism (MP) methods and automatic parallelism (AP) methods. In MP methods  [37, 30, 44], one or several parallel strategies are manually optimized by researchers or developers. MP methods require extensive domain knowledge in deep learning models and hardware architectures. With the rapid development of deep learning models and the increasing diversity of modern hardware architectures [11, 12], MP methods demand considerable human effort and have limited flexibility.
To address the two limitations of MP methods, AP methods [28, 14, 47] have recently been proposed for automating the parallel strategy optimization process. Although existing AP methods have achieved promising progress, they optimize the two categories of parallel strategies separately rather than jointly. More specifically, some methods optimize only one category of parallel strategies [16, 42, 17, 36, 46, 2, 23], and the others optimize inter- and intra-layer parallelism hierarchically [28, 38, 39, 9, 14, 47]. Hence, existing AP methods suffer from sub-optimal solutions.
@@ 2.2 Parallel method
(a) Model
(b) MP
(c) Inter-layer-only AP
(d) Intra-layer-only AP
(e) Hierarchical AP
(f) UniAP
Figure 1: Parallel methods for optimizing parallel strategies for a three-layer model in (a). Numbered boxes represent the corresponding layers of the model. Layers with gray backgrounds represent the final parallel strategy for each method. Layers without dotted lines indicate that DP was applied to them, while TP was applied to the others. Solid arrows between layers represent connections within a single pipeline stage, while dashed arrows represent connections between consecutive pipeline stages.
@@ 3 Method
In this section, we introduce our proposed method called UniAP, which jointly optimizes the two categories of parallel strategies, including PP, DP, TP, and FSDP, to find an optimal solution. Figure 1 illustrates the difference between UniAP and other parallel methods. MP methods are manually designed with limited flexibility. Inter-layer-only and intra-layer-only AP methods optimize (search) from a set of candidate inter-layer-only and intra-layer-only parallel strategies, respectively. Hierarchical AP methods first adopt greedy or dynamic programming to propose candidate inter-layer parallel strategies. Then, they optimize the intra-layer parallel strategy for every fixed inter-layer parallel strategy. UniAP has the largest strategy space for exploration (joint optimization).
@@ 3.1 Profiling
UniAP collects runtime information about the hardware environment and deep learning model during profiling. For the hardware environment, UniAP evaluates the efficiency of all-reduce and point-to-point (P2P) communication for different device subsets. For example, when profiling a node with 4 GPUs, UniAP measures the all-reduce efficiency for various DP, TP, and FSDP combinations across these GPUs. Additionally, UniAP ranks these GPUs from 0 to 3 and evaluates the speed of P2P for two pipeline options: ($0\rightarrow 2$ and $1\rightarrow 3$) and ($0\rightarrow 1$, $1\rightarrow 2$ and $2\rightarrow 3$). Furthermore, UniAP estimates the computation-communication overlap coefficient (CCOC) [25, 33].
@@ 3.2 Cost model
Time cost model To estimate computation time, UniAP first calculates the forward computation time by multiplying the batch size with the forward computation time per sample obtained from profiling. Users can obtain a more precise result by profiling the forward time on the specified batch size. For Transformer-based models that mainly consist of the MatMul operator, the computation time in the BP stages is roughly twice that of the FP stages [30, 21, 25]. Additionally, UniAP estimates the communication time by dividing the size of transmitting tensors by the profiled communication efficiency for different communication primitives. To accommodate overlapping, UniAP multiplies the profiled CCOC by the overlapping interval of computation and communication. To model the communication time between pipeline stages, UniAP calculates the cross-stage cost between consecutive stages by the summation of P2P costs.
@@
For example, if we choose precision as FP32, then $c_{dtype}=(4+4+4+4)/4=4$ since the learnable parameters, gradients, momentum and variance consume equal size of memory. If we opt for mixed precision with FP16 activated, then $c_{dtype}=(4+4+4+2+2)/2=8$. Finally, UniAP aggregates the activation memory cost $m_{a}$, memory cost of model states $m_{s}$, and context memory cost $m_{c}$ to a constant matrix M, where $\textbf{{M}}_{uk}$ denotes the memory cost for the $k$-th intra-layer strategy of layer $u$ on a single device.
@@ 3.3.1 Objective function
The objective function tries to minimize TPI. In this paper, we have chosen GPipe as our PP strategy for illustration. [^2] UniAP is also compatible with other PP strategies. For example, users need to modify only the memory constraint in Section 3.3.2 to adapt to synchronous 1F1B pipeline [29, 9]. Figure 4 depicts the time cost decomposition of a GPipe-style PP with non-negligible communication costs. The time needed to apply gradients at the end of each iteration is not included, as it depends on the optimizer and is insignificant compared to the total time spent on FP and BP.
We denote the cost for computation stages as =$\mathbb{P}=\{p_{1},p_{2},\dots,p_{deg}\}$ and the cost for communication stages as $\mathbb{O}=\{o_{1},o_{2},\dots,o_{deg-1}\}$. Here, $deg$ represents the number of computation stages, which corresponds to the pipeline parallel size. In Figure 4, $fp_{i}$ and $bp_{i}$ denote forward and backward computation time for computation stage $i$, respectively. $fo_{j}$ and $bo_{j}$ denote forward and backward communication time for communication stage $j$, respectively. Hence, we have $p_{i}=fp_{i}+bp_{i}$ and $o_{j}=fo_{j}+bo_{j}$.
@@
$\min tpi_{gpipe}$ $=\sum_{i=1}^{deg}p_{i}+\sum_{j=1}^{deg-1}o_{j}$ (2)
$+(c-1)\max\left(\mathbb{P}\cup\mathbb{O}\right).$
@@ 3.3.2 Constraint
We first introduce additional notations before presenting the constraints. For a given layer $u\in\mathbb{V}$, $\mathbb{S}_{u}$ represents its set of intra-layer parallel strategies, $\textbf{{A}}_{uk}$ denotes the $k$-th intra-layer execution cost obtained from our time cost model. Additionally, we use $\textbf{{S}}_{uk}\in\{0,1\}$ to indicate whether the $k$-th parallel strategy is selected for the layer $u$, and use $\textbf{{P}}_{ui}\in\{0,1\}$ to indicate whether layer $u$ is placed on the $i$-th computation stage. Each edge $\langle u,v\rangle\in\mathbb{E}$ is assigned a resharding cost denoted by $\textbf{{R}}_{uv}$ if the vertices are located within the same pipeline stage. Alternatively, if the vertices are located across consecutive stages, the resharding cost between them is denoted by $\textbf{{R}}^{\prime}_{uv}$. These two resharding costs are constant matrices derived from our time cost model.
@@
$\sum_{u\in\mathbb{V}}\textbf{{P}}_{ui}\textbf{{S}}_{u}^{\mathsf{T}}\textbf{{A}}_{u}+\sum_{\langle u,v\rangle\in\mathbb{E}}\textbf{{P}}_{ui}\textbf{{P}}_{vi}(\textbf{{S}}_{u}^{\mathsf{T}}\textbf{{R}}_{uv}\textbf{{S}}_{v})=p_{i},$ (3)
$~{}\forall i\in\{1,\dots,deg\}.$
@@
$\sum_{\langle u,v\rangle\in\mathbb{E}}\textbf{{P}}_{uj}\textbf{{P}}_{v(j+1)}(\textbf{{S}}_{u}^{\mathsf{T}}\textbf{{R}}^{\prime}_{uv}\textbf{{S}}_{v})=o_{j},$ (4)
$~{}\forall j\in\{1,\dots,deg-1\}.$
@@
$\sum_{u\in\mathbb{V}}\textbf{{P}}_{ui}\textbf{{S}}_{u}^{\mathsf{T}}\textbf{{M}}_{u}\leqslant m,~{}\forall i\in\{1,\dots,deg\}.$ (5)
@@
Order-preserving constraint PP is not a single-program multiple-data (SPMD) parallel strategy [15]. Hence, we need an order-preserving constraint to ensure that the subgraphs of $\mathcal{G}$ are contiguous. We adopt the definition of contiguous from previous work [38, 39].
@@
$\textbf{{Z}}_{vi}\geqslant\textbf{{P}}_{vi},$
$~{}\forall v\in\mathbb{V},~{}\forall i\in\{1,2,\dots,deg\},$ (6a)
$\textbf{{Z}}_{vi}\leqslant\textbf{{Z}}_{ui},$
$~{}\forall u,v\in\mathbb{V},~{}\forall\langle u,v\rangle\in\mathbb{E},~{}\forall i\in\{1,2,\dots,deg\},$ (6b)
$\textbf{{Z}}_{vi}\leqslant\textbf{{P}}_{vi}-\textbf{{P}}_{ui}+1,$
$~{}\forall u,v\in\mathbb{V},~{}\forall\langle u,v\rangle\in\mathbb{E},~{}\forall i\in\{1,2,\dots,deg\}.$ (6c)
Detailed proof can be found in Appendix A.
@@
$\sum_{i=1}^{deg}\textbf{{P}}_{ui}=1,$ $\forall u\in\mathbb{V},$ (7a)
$\sum_{u\in\mathbb{V}}\textbf{{P}}_{ui}\geqslant 1,$ $\forall i\in\{1,\dots,deg\},$ (7b)
$\textbf{{P}}_{ui}\in\{0,1\},$ $\forall u\in\mathbb{V},~{}i\in\{1,\dots,deg\}.$ (7c)
@@
$\sum_{k=1}^{\lvert\mathbb{S}_{u}\rvert}\textbf{{S}}_{uk}=1,$ $\forall u\in\mathbb{V},$ (8a)
$\textbf{{S}}_{uk}\in\{0,1\},$ $\forall u\in\mathbb{V},~{}k\in\{1,\dots,|\mathbb{S}_{u}|\}.$ (8b)
@@
  Output: Optimal cost $cost^{*}$, pipeline parallel size $deg^{*}$, the number of micro-batches $c^{*}$, layer placement $\textbf{{P}}^{*}$, and intra-layer strategy $\textbf{{S}}^{*}$.
  $deg^{*}=1$;
@@
  $cost^{*}$, $P^{*}$, $S^{*}$ = QIP(A, R, M);
@@
        A, R, $\textbf{{R}}^{\prime}$, M = CostModeling($PR$, $SD[deg]$, $\mathcal{G}$, $b$);
        $cost$, P, S = MIQP(A, R, $\textbf{{R}}^{\prime}$, M, $deg$, $c$);
@@
           $cost^{*}$, $deg^{*}$, $c^{*}$, $\textbf{{P}}^{*}$, $\textbf{{S}}^{*}$ = $cost$, $deg$, $c$, P, S;
@@
3.4 Unified optimization process
UOP integrates the cost model and MIQP based on the profiling results and the computation graph to return the optimal parallel strategy and the corresponding TPI. Algorithm 1 summarizes the whole process. In Algorithm 1, we denote intra-layer cost as A, inter-layer cost as R, cross-stage cost as $\textbf{{R}}^{\prime}$, and memory cost as M. The CostModeling process calculates these four costs based on the cost model described in Section 3.2.
First, UOP optimizes intra-layer-only parallelism for cases in which pipeline parallelism is not adopted. Several works [47, 23] have used quadratic integer programming (QIP) to optimize intra-layer-only parallel strategy and achieved promising results. UniAP provides a QIP formulation for intra-layer-only parallelism in Appendix B.
Then, UOP enumerates all factors of $n$ except 1 as the pipeline parallel size $deg$. For each $deg$, UOP enumerates all factors of $B$ except 1 as the number of micro-batches $c$. These enumerations aim to achieve load balance on a homogeneous cluster.
For each candidate $deg$ and $c$, UOP formulates the cost for a training iteration to an MIQP expression. It then waits for the MIQP solver to return the optimal cost and parallel strategy. Specifically, our implementation adopts the Gurobi Optimizer [13] as our QIP and MIQP solver.
Finally, UOP returns the minimum cost $cost^{*}$ and its corresponding pipeline parallel size $deg^{*}$, number of micro-batches $c^{*}$, layer placement $\textbf{{P}}^{*}$, and intra-layer strategies $\textbf{{S}}^{*}$. We provide visualization for a candidate solution to UOP in Appendix C.